For, A→ R, liquid phase reaction carried out in a 1000 liter MFR with CAo=10 mol/lit, v=10 lit/min, -rA = 0.9 mol .lit¹.min¹ ,what is the exit concentration of A (in mol/liter)? ( )





A)1




B)3




C)5




D)7


Solve with steps ?

1
Expert's answer
2022-02-18T04:51:03-0500

t = V/v = 1000 l / 10 l/min = 100 min

CAr = 0.09 mol/l min x 100 min = 9 mol/l

CA = 10 mol/l - 9 mol/l = 1 mol/l

A) 1